identify the type of the differential equation y =xdy/dx+1-logdy/dx and hance solve it
1
Expert's answer
2020-05-18T16:51:12-0400

We have given the differential equation as

"y =x\frac{dy}{dx}+1-log\big(\frac{dy}{dx}\big) \hspace{1cm}(\clubs)"

Clearly, it has the form of

"y =x\big(\frac{dy}{dx}\big)+1-H\big(\frac{dy}{dx}\big)"

i.e

"H=H(\frac{dy}{dx})"

Thus, this form of the differential equation is called Clairaut's equation.

Let's fix the notation for

"y'=\frac{dy}{dx} \& \: y''=\frac{d^2y}{dx^2}"

Now, to find the solution of the equation "(\clubs)" ,differentiate both side of the equation with respect to "x" ,hence we get,


"y'=xy''+y'-\bigg(\frac{1}{y'}y''\bigg)\\
\implies \big(x-\frac{1}{y'}\big)y''=0 \hspace{1cm}(\spades)"

Clearly, equation "(\spades)" has two cases.

Case-I:

Suppose

"y''=0"

we already know the solution(which is trivial) i.e

"y'=a \\
\implies y=ax+b"

where,"a,b" are any constant value.

Now, if we plugin the above solution to "(\clubs)" we get,


"y=ax+1-log(a) \iff y=ax+b"

Case-II:

Suppose

"x-\frac{1}{y'}=0\\
\implies y'=\frac{1}{x}\\
\implies \int dy=\int\frac{1}{x}dx\\
\implies y=log(x)+c"

If we plugin the above solution to "(\clubs)" then we get,

"c=2"

Hence, the final solution will be in the form of

"y=log(x)+2"

Interestingly, this solution is singular!

Hence, we are done.